OpenStudy (anonymous):

The sum of one-third of a number and 5 is 27. Find the number.

OpenStudy (anonymous):

I believe this equates to 1/3x+5=27

OpenStudy (anonymous):

i do too

OpenStudy (anonymous):

is the answer 76?

OpenStudy (shamim):

x=66?

OpenStudy (anonymous):

how did you get that?

sammixboo (sammixboo):

To check if it is 76, plug it in where the variable it

sammixboo (sammixboo):

\(\dfrac{1}{3}\times76+5=27\)

OpenStudy (shamim):

x/3=27-5 x/3=22 x=22*3=?

OpenStudy (anonymous):

oh i have got it now :) thank you very much

sammixboo (sammixboo):

=)